OpenStudy (anonymous):

multiple algebra 2 questions. please help :)

OpenStudy (anonymous):

what are the questions?

OpenStudy (anonymous):

2 divided by 6- sqrt of 3. answer choices. 12+ 2 sqrt of 3 over 33, 12+ sqrt of 3 over 33, 2 sqrt of 3 over 6 sqrt of 3 - 3, 2 sqrt of 3 over 3.

OpenStudy (anonymous):

in your question is there a -ve sign after the 6 or is that an error?

OpenStudy (anonymous):

nope, just 2 over 6 minus sqrt of 3

OpenStudy (anonymous):

\[2/(6-\sqrt{3}) = 2(6+\sqrt{3)}/36-3 = 2(6+\sqrt{3)}/33\] \[=(12+2\sqrt{3})\] I hope you got that

OpenStudy (anonymous):

\[=(12+2\sqrt{3})/33\]

OpenStudy (anonymous):

wow thankyou so much :) @gitahimart